OpenStudy (anonymous):

2x^2-16x+9=0 solve by completing the square

jimthompson5910 (jim_thompson5910):

Answers are \[\large x=\frac{8+\sqrt{46}}{2} \ \textrm{ or } \ \ x=\frac{8-\sqrt{46}}{2}\] See attached image for how I got those answers.